OpenStudy (anonymous):

sqrt 100x=x+16 Help me solve this please? The 100x is covered in square root symbol, its a quadratic

OpenStudy (anonymous):

Raise both sides to the second power to get rid of the root. Then solve as a usual quadratic equation.

OpenStudy (anonymous):

So that, for your convenience, 100x = (x+16)^2 100x = x^2+32x+256 x^2-68x+256=0

OpenStudy (anonymous):

you could square both sides, and then foil out the right side. and on the left side, since you squared it, the square out sign cancels out [i.e. recall that squareroot is the the inverse of squaring.]

OpenStudy (anonymous):

x= 4 and x= 64